OpenStudy (anonymous):

Solve the Equation 4d+2d-4d+4=3d A.) 4/7 B.) 4 C.) -4/5 D.) -4

OpenStudy (anonymous):

b) 4

OpenStudy (anonymous):

:( you shouldn't give the answer. she is perfectly capable of learning how to solve this equation.

OpenStudy (anonymous):

dont even -_-

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

I dont see the point in simply giving her the answer, this is about learning how to do different math problems. If she doesn't understand the process, shes learned nothing.

OpenStudy (anonymous):

Just sum the equals and then pass the number for the another side of the equation.

OpenStudy (anonymous):

\[4d+2d-4d+4=3d \] \[2d+4=3d\] \[2d(-2d)+4=3d(-2d)\] \[4=d \]